OpenStudy (anonymous):

Solve equation by taking square roots 8a^2 -7= -56

OpenStudy (cwrw238):

add 7 to both sides 8a^2 = -49 a^2 = -49 / 8 hmm - can you see a problem here ? have you studied complex numbers?

OpenStudy (cwrw238):

there are no real roots because the right hand side is negative Are you sure you have written the question correctly?

OpenStudy (anonymous):

the question is typed correctly

OpenStudy (anonymous):

the answer has imaginary numbers and has extraneous solutions

OpenStudy (cwrw238):

ok taking square roots:- a = 7i / sqrt8 or -7i / sqrt8 can u continue from here?

OpenStudy (anonymous):

yes thank you

OpenStudy (cwrw238):

yw